OpenStudy (anonymous):

show all steps,,,,2(n - 3) + 5 = 3(n - 1)

OpenStudy (anonymous):

2n-6+5=3n-3 isolate variable n=2

OpenStudy (anonymous):

2n-6 + 5 = 3n-3 n = -6+5+3 n = 2